Use the formula N=Ie^kt, where N is the number of items of the initial population "I", at the time "t", and "k" is the growth constant equal to the percent of growth per unit of time. A certain radioactive isotope has a half-life of approximately 1750 years. How many years would be required for a given amount of this isotope to decay to 25% of that amount?
Answer by Boreal(15235)

This would be 2 half-lives or 3500 years
N=Ie^(1750k)
divide by I and N/I=1/4
1/2=e^(1750k)
ln of both sides
-ln2=1750k
k=-ln2/1750=-0.000396
N/I=0.25=e^(0.000396*t)
ln(0.25)=0.000396t
-ln4/-0.000396=t=3500.74 years, but 3500 is exact.